A rectangular game board's long side is 2 in. longer than the short side. The area of the board is 288 in^2
Find the dimensions of the game board.

Answer by JulietG(1812)

Here's one way to do it:
Area = length x width
You're given the area and the sides
288 = (x+2) * x
288 = x^2 + 2x
x^2 + 2x - 288 = 0
Factor
x = -18, 16
Ignore the negative since the measurement can't be negative.
The sides are 18 length and 16 width